OpenStudy (anonymous):

what is the simplified form of x-5/x^2-3x-10*x+2/x^2+x-12

OpenStudy (anonymous):

if this is the problem, factoring should do the trick: \(\large \frac{x-5}{x^2-3x-10} \cdot \frac{x+2}{x^2+x-12} \) can you give me the factors of each denominator?

OpenStudy (anonymous):

would it be 1/(x+2)(x-5)?